# Causal Explanations for Image Classifiers

### v2

@@ 1 Introduction
RQ3
Is there an efficiently computable approximation to explanations?
RQ4
What are suitable quality measures for explanations? A measure is suitable if it is meaningful with respect to quality and produces results that match our intuition.
RQ5
What is the quality of explanations computed by our algorithms compared to other black-box methods?
RQ6
RQ7
[…]
We compare the experimental results of our implementation of the algorithm in a tool rex with those of other black-box tools on a number of benchmark datasets: ImageNet, VOC2012, and a set of partially occluded images. Our results show that rex is on par with the most efficient tools in terms of execution time, while producing the smallest explanations. We also analyse the explanations using the standard measures of insertion and deletion curves and the intersection with the background or the occlusion. The results show that rex outperforms the other tools on insertion curves and the overlap with irrelevant parts of the image. We discuss the logic behind deletion curves and demonstrate that low deletion is a measure of the quality of the model for particular types of images, rather than for the quality of an explanation. We also argue that explanations computed by rex are so small that any further reduction in their size is meaningless.
The rest of the paper is organized as follows. Section 2 presents an overview of the related work on black-box explainable AI. Section 3 gives the necessary background on actual causality. Section 4 provides the theoretical foundations of our approach to explainability. Section 5 is a detailed overview of our algorithm, with the evaluation results discussed in Section 6. The tool rex is open source and available at https://github.com/ReX-XAI/ReX. All benchmark sets are publicly available and referenced in Section 6.
2 Related Work
There is a large body of work on explaining the results of image classifiers. As our primary interest is in black-box approaches, we only give a brief overview of the white-box algorithms here. The existing approaches to post hoc explainability can be largely grouped into two categories: propagation and perturbation.
Propagation-based explanation methods are often regarded as more efficient. They back-propagate a model’s decision to the input layer to determine the weight of each input feature for the decision. Grad-CAM by (?) only needs one backward pass and propagates the class-specific gradient into the final convolutional layer of a DNN to coarsely highlight important regions of an input image. Guided Back-Propagation (gbp) by (?) computes the single and average partial derivatives of the output to attribute the prediction of a DNN. Integrated Gradients (ig) by (?) further uses two axioms called sensitivity and completeness for the problem of how to attribute the classification by a deep network to its input features. noisetunnel by (?) adds gaussian noise to the image before applying another XAI method. The noise serves to smooth any resultant saliency map, improving explanation quality. shap (SHapley Additive exPlanations) by (?) goes beyond propagation and identifies inputs that are similar to the input for which the output is to be explained, and ranks the features of the input according to their difference. A key advantage of shap is that it does not require counterfactuals. gradientshap is gradient-based method that approximates Shapley values via gradient integration and a noise tunnel. Both ig and gradientshap require access to model gradients while computing explanations. They form part of a family of whitebox explanation methods which differ from the Grad-CAM family in that they do not rely on examination of specified activation layers in the model. Other methods for approximating Shapley values are numerous, among which KernelShap (?) and DeepLiftShap (?) are probably the most used. KernelShap uses the lime framework to approximate Shapley values, whereas DeepListShap uses the DeepLift (?) method to the same end.
In contrast to propagation-based explanation methods, perturbation-based explanation approaches explore the input space directly in search for an explanation. The exploration/search often requires a large number of inference passes, which incurs significant computational cost when compared to propagation methods. There are many sampling methods, most of which are based on random search or heuristics. The chief advantage of perturbation-based methods is that they are model independent and do not require any access to model internals. Given a particular input, lime by (?) samples the the neighborhood of this input and creates a inherently explainable model to approximate the system’s local behavior; owing to the high computational cost of this approach, the ranking uses super-pixels via segmentation instead of individual pixels. (?) replace the natural distribution of the input by a user-defined distribution and then use the Shapley Value method to analyze combinations of input features and to rank their importance. (?) estimate the importance of input features by measuring the the flow of information between inputs and outputs. Both the Shapley Value and the information-theoretic approaches are computationally expensive. In rise by (?), the importance of a pixel is computed as the expectation over all local perturbations conditioned on the event that the pixel is observed. (?) have introduced the concept of “extreme perturbations” to improve the perturbation analysis in the extremal algorithm. More recently, spectrum-based fault localisation (SBFL) has been applied to explaining image classifiers. The technique has been implemented in the tool deepcover by (?).
The method and tool, rex, presented in this paper is a perturbation-based approach and addresses the limitations of existing black-box methods in two aspects. The feature masking in rex uses causal reasoning that provides a guarantee (subject to the assumption). Owing to its guided iterative refinement, rex is computationally efficient, while still producing explanations that are the state of the art.
3 Background on Actual Causality
In this section, we briefly review the definitions of causality and causal models introduced by (?) and relevant definitions of causes and explanations in image classification by (?). The reader is referred to (?) for further reading.
Formally, a causal model $M$ is a pair $({\mathcal{}S},{\mathcal{}F})$, where ${\mathcal{}S}$ is a signature, which explicitly lists the endogenous and exogenous variables and characterizes their possible values, and ${\mathcal{}F}$ defines a set of (modifiable) structural equations, relating the values of the variables. A signature ${\mathcal{}S}$ is a tuple $({\mathcal{}U},{\mathcal{}V},{\mathcal{}R})$, where ${\mathcal{}U}$ is a set of exogenous variables, ${\mathcal{}V}$ is a set of endogenous variables, and ${\mathcal{}R}$ associates with every variable $Y\in{\mathcal{}U}\cup{\mathcal{}V}$ a nonempty set ${\mathcal{}R}(Y)$ of possible values for $Y$ (i.e., the set of values over which $Y$ ranges). For simplicity, we assume here that ${\mathcal{}V}$ is finite, as is ${\mathcal{}R}(Y)$ for every endogenous variable $Y\in{\mathcal{}V}$. The set ${\mathcal{}F}$ associates with each endogenous variable $X\in{\mathcal{}V}$ a function denoted $F_{X}$ (i.e., $F_{X}={\mathcal{}F}(X)$) such that $F_{X}:(\times_{U\in{\mathcal{}U}}{\mathcal{}R}(U))\times(\times_{Y\in{\mathcal{}V}-\{X\}}{\mathcal{}R}(Y))\rightarrow{\mathcal{}R}(X)$.
Probabilistic causal models are pairs $(M,\Pr)$, where $M$ is a causal model and $\Pr$ is a probability on the contexts. A causal model $M$ is recursive (or acyclic) if its causal graph is acyclic. If $M$ is an acyclic causal model, then given a context, that is, a setting $\vec{u}$ for the exogenous variables in ${\mathcal{}U}$, the values of all the other variables are determined. In this paper we restrict the discussion to recursive models.
We call a pair $(M,\vec{u})$ consisting of a causal model $M$ and a context $\vec{u}$ a (causal) setting. A causal formula $\psi$ is true or false in a setting. We write $(M,\vec{u})\models\psi$ if the causal formula $\psi$ is true in the setting $(M,\vec{u})$. The $\models$ relation is defined inductively. $(M,\vec{u})\models X=x$ if the variable $X$ has value $x$ in the unique solution to the equations in $M$ in context $\vec{u}$. Finally, $(M,\vec{u})\models[\vec{Y}\leftarrow\vec{y}]\varphi$ if $(M_{\vec{Y}=\vec{y}},\vec{u})\models\varphi$, where $M_{\vec{Y}\leftarrow\vec{y}}$ is the causal model that is identical to $M$, except that the variables in $\vec{Y}$ are set to $Y=y$ for each $Y\in\vec{Y}$ and its corresponding value $y\in\vec{y}$.
A standard use of causal models is to define actual causation: that is, what it means for some particular event that occurred to cause another particular event. There have been a number of definitions of actual causation given for acyclic models (e.g., (?????????)). In this paper, we focus on what has become known as the modified Halpern–Pearl definition and some related definitions introduced by (?). We briefly review the relevant definitions below. The events that can be causes are arbitrary conjunctions of primitive events (formulas of the form $X=x$); the events that can be caused are primitive events, denoting the output of the model.
Definition 3.1 (Actual cause).
$\vec{X}=\vec{x}$ is an actual cause of $\varphi$ in $(M,\vec{u})$ if the following three conditions hold:
AC1.
$(M,\vec{u})\models(\vec{X}=\vec{x})$ and $(M,\vec{u})\models\varphi$.
AC2.
AC3.
In the special case that $\vec{W}=\emptyset$, we get the but-for definition. A variable $x$ in an actual cause $\vec{X}$ is called a part of a cause. In what follows, we adopt the convention of Halpern and state that part of a cause is a cause.
The notion of explanation taken from (?) is relative to a set of contexts.
Definition 3.2 (Explanation).
$\vec{X}=\vec{x}$ is an explanation of $\varphi$ relative to a set ${\mathcal{}K}$ of contexts in a causal model $M$ if the following conditions hold:
EX1a.
If $\vec{u}\in{\mathcal{}K}$ and $(M,\vec{u})\models(\vec{X}=\vec{x})\wedge\varphi$, then there exists a conjunct $X=x$ of $\vec{X}=\vec{x}$ and a (possibly empty) conjunction $\vec{Y}=\vec{y}$ such that $X=x\wedge\vec{Y}=\vec{y}$ is an actual cause of $\varphi$ in $(M,\vec{u})$.
EX1b.
$(M,\vec{u}^{\prime})\models[\vec{X}=\vec{x}]\varphi$ for all contexts $\vec{u}^{\prime}\in{\mathcal{}K}$.
EX2.
$\vec{X}$ is minimal; there is no strict subset $\vec{X}^{\prime}$ of $\vec{X}$ such that $\vec{X}^{\prime}=\vec{x}^{\prime}$ satisfies EX1, where $\vec{x}^{\prime}$ is the restriction of $\vec{x}$ to the variables in $\vec{X}^{\prime}$. (This is SC4).
EX3.
$(M,u)\models\vec{X}=\vec{x}\wedge\varphi$ for some $u\in{\mathcal{}K}$.
4 Theoretical Foundations
We approach the first two research questions theoretically.
RQ1
We view an image classifier (e.g. a neural network) as a probabilistic causal model. Specifically, the endogenous variables are taken to be the set $\vec{V}$ of pixels that the image classifier gets as input, together with an output variable that we call $O$. The variable $V_{i}\in\vec{V}$ describes the color and intensity of pixel $i$; its value is determined by the exogenous variables. The equation for $O$ determines the output of the model as a function of the pixel values. Thus, the causal network has depth 222, with the exogenous variables determining the feature variables, and the feature variables determining the output variable.
We also assume causal independence between the feature variables $\vec{V}$. In our setting, $\vec{V}$ is the set of pixels of the input image, with the values capturing the color and intensity of the pixels. Pixel independence is a common assumption in explainability tools. This is a non-trivial assumption, and it might seem far-fetched, especially if we consider images capturing real objects: if a group of pixels captures, say, a cat’s ear, then a group of pixels below it should capture a cat’s eye. However, we argue that it is, in fact, accurate on images. Indeed, consider the “Photobombing” dataset, described in Section 6.1. This dataset is comprised of partially obscured images, obtained by overlaying random color patches over Imagenet images. An input from this dataset is a perfectly valid image. Indeed, obscuring a part of the input image either by introducing an artificial object or by positioning a real object in front of the primary subject of the classification does not lead to any change in unobscured pixels. For example, obscuring a cat’s ear does not lead to any change in an (unobscured) group of pixels below it, capturing a cat’s eye. Thus, pixel independence holds on general images.
We note that pixels are not concepts: a  [^2] D image is a projection of a scene in the  [^3] D real world on a plane; concepts that are present in the object in a real world can be invisible on this projection, hence the pixel independence.
We refer the reader to (?) for a more in-depth discussion of causal independence between pixel values in image classification. We note that the causal independence assumption is not true in other types of inputs, such as tabular or spectral data. For those types of inputs, assuming independence is clearly an approximation and might lead to inaccurate results. In this paper, however, we focus on images, where causal independence between pixels holds, and it greatly simplifies explanation extraction, as we show below.
Moreover, as the causal network is of depth 222, all parents of the output variable $O$ are contained in $\vec{V}$. Given these assumptions, the probability on contexts directly corresponds to the probability on seeing various images (which the model presumably learns during training).
Given an input image $I$, the set of contexts ${\mathcal{}K}$ that we consider for an explanation is the set ${\mathcal{}K}_{I}$ obtained by all partial occlusions of $I$, where an occlusion sets a part $\vec{Y}$ of the image to a predefined masking color $\vec{y}$. The probability distribution over ${\mathcal{}K}_{I}$ is assumed to be uniform.
Under the assumptions above, the following definition is equivalent to Definition 3.2, as shown by (?).
Definition 4.1 (Explanation for image classification).
An explanation is a minimal subset of pixels of a given input image that is sufficient for the model $\mathcal{N}$ to classify the image, where “sufficient” is defined as containing only this subset of pixels from the original image, with the other pixels set to the masking color.
As our goal is ranking of the pixels according to their importance for the classification, we only consider singleton causes, that is, causes consisting of one pixel and its value. Moreover, for each pixel we only consider two possible values: its original value in the image and a masking color.
Definition 4.2 (Singleton cause for image classification).
For an image $x$ classified by the DNN as $f(x)=o$, a pixel $p_{i}$ of $x$ and its value in $x$ is a cause of $o$ iff there exists a subset $P_{j}$ of pixels of $x$ such that the following conditions hold:
SC1.
$p_{i}\not\in P_{j}$;
SC2.
changing the color of any subset $P^{\prime}_{j}\subseteq P_{j}$ to the masking color does not change the classification;
SC3.
changing the color of $P_{j}$ and the color of $p_{i}$ to the masking color changes the classification.
We call such $P_{j}$ a witness to the fact that $p_{i}$ is a cause of $x$ being classified as $o$.
Definition 4.2 is equivalent to a part of a cause according to Definition 3.1 in binary causal models with all variables being independent of each other. Moreover, under these restrictions, Definition 4.2 is equivalent to the following definition, which is the original definition of an actual cause by (?).
Definition 4.3 (Original definition of actual cause).
$\vec{X}=\vec{x}$ is an actual cause of $\varphi$ in $(M,\vec{u})$ if the following three conditions hold:
AC1.
$(M,\vec{u})\models(\vec{X}=\vec{x})$ and $(M,\vec{u})\models\varphi$.
AC2’.
There exists a partition $(\vec{Z},\vec{W})$ of $\vec{V}$ with $\vec{X}\subseteq\vec{Z}$ and some setting $(\vec{x}^{\prime},\vec{w}^{\prime})$ of the variables in $(\vec{Z},\vec{W})$ such that if $(M,\vec{u})\models Z=z*$ for $Z\in\vec{Z}$, then
  1. 1.
    $(M,\vec{u})\models[\vec{X}\leftarrow\vec{x}^{\prime},\vec{W}\leftarrow\vec{w}^{\prime}]\neg\varphi$. In other words, changing the values of $(\vec{X},\vec{W})$ from $(\vec{x},\vec{w})$ to $(\vec{x}^{\prime},\vec{w}^{\prime})$ changes $\varphi$ from true to false;
  2. 2.
    $(M,\vec{u})\models[\vec{X}\leftarrow\vec{x},\vec{W}\leftarrow\vec{w}^{\prime},\vec{Z}^{\prime}\leftarrow\vec{z}*]\varphi$ for all subsets $\vec{Z}^{\prime}$ of $\vec{Z}$. In other words, setting $\vec{W}$ to $\vec{w}^{\prime}$ should have no effect on $\varphi$ as long as $\vec{X}$ is kept at its current value $\vec{x}$, even if all the variables in an arbitrary subset of $\vec{Z}$ are set to their original values in the context $\vec{u}$.
AC3.
$\vec{X}$ is minimal; there is no strict subset $\vec{X}^{\prime}$ of $\vec{X}$ that satisfies AC1 and AC2’.
The equivalence is based on the following observations. The minimality is satisfied due to only considering singletons. The set $\vec{Z}$ does not play a role due to the independence between the pixels. Removing $\vec{Z}$ and the minimality condition from Definition 4.3 we get Definition 4.2.
Computing causality in image classification is, therefore, NP-complete, based on the above equivalence and the result by (?), who proved NP-completeness of Definition 4.3 for binary causal models. Moreover, following the same reasoning, Definition 4.1 is equivalent to the definition of explanation by (?) for binary causal models, which is DP-complete (?). Hence, there is little hope to find an efficient algorithm for computing exact explanations for image classification.
Finally, the following definition of the degree of responsibility holds for the case of image classifiers and singleton causes as in Definition 4.2.
Definition 4.4 (Simplified responsibility).
[…]
RQ2
[…]
Observation 4.5.
Given an image $x$ and its classification $o$, we can calculate the degree of responsibility of each pixel $p_{i}$ of $x$ by directly applying Definition 4.2, that is, by checking the conditions SC1, SC2, and SC3 for all subsets $P_{j}$ of pixels of $x$ and then choosing a smallest witness subset. While there is an underlying Boolean formula that determines the classification $o$ given the values of the pixels of $x$, we do not need to discover this formula in order to calculate the degree of responsibility of each pixel of $x$.
In the next section we introduce a compositional approach to computing approximate explanations. The main idea is first to rank pixels according to their approximate degree of responsibility, and then use a greedy approach to extract (approximate) explanations. The algorithm for computing an approximate degree of responsibility is based on the notion of a super-pixel $P_{i}$, which is a subset of pixels of a given image. Given an image $x$, we partition it into a small number of superpixels and compute their degree of responsibility for the output of the DNN. Then, we only refine those superpixels with a high responsibility (exceeding a predefined threshold). The scalability of the approach relies on the following observation, which is shown to be heuristically true in our experiments.
Observation 4.6.
The pixels with the highest responsibility for the DNN’s decision are located in super-pixels with the highest responsibility.
Intuitively, the observation holds when pixels with high responsibility do not appear in the superpixels surrounded by other pixels with very low responsibility for the input image classification outcome. While this can happen in principle, we do not encounter this case in practice owing to the continuous nature of images (even when the explanation is non-contiguous). This property is key to the success of our algorithm.
5 Compositional Explanations
@@
Figure 2: High level overview of Algorithm 3. Iterating parts 1 and 2 of the algorithm produce a responsibility map (①). The pixels in the image are then ordered by their approximate responsibility and used to produce a minimal sufficient explanation (②), which captures all that the DNN needs to give a classification 'starfish'.
We recall the third research question.
RQ3
Is there an efficiently computable approximation to explanations?
We answer this question affirmatively, by presenting our greedy compositional explanation (CE) algorithm. The general idea is to calculate the responsibility of a superpixel and recursively distribute this responsibility to all pixels within this superpixel. The CE approach in this work consists of the following steps.
  1. 1.
    Given a set of superpixels, compute the responsibility of each superpixel (Section 5.1).
  2. 2.
    Following the responsibility result in Step 1, further refine the superpixel and calculate the responsibility for the refined superpixels (Section 5.2).
@@
We note that the ranking depends on the selected partition. To ameliorate the effect of a particular partition, the algorithm is re-run a number of times with partitions selected independently at random, and the results are averaged across all the partitions (see Section 5.3). Section 5.4 gives a step-by-step example to illustrate the working of the algorithm.
5.1 Computing the responsibility of a superpixel
@@
Given a set of pixels $\mathcal{P}$, we use $\mathds{P}_{i}$ to denote a partition of $\mathcal{P}$, that is, a set $\{P_{i,j}:\bigcup P_{i,j}=\mathcal{P}$ and $\forall j\not=k,P_{i,j}\cap P_{i,k}=\emptyset\}$. The number of elements in $\mathds{P}_{i}$ is a parameter, denoted by $s$; in this work, we consider $s=4$. We refer to $P_{i,j}$ as superpixels. It is insufficient to consider each superpixel in isolation: there is no reason why anyone super pixel should be the cause of a classification. Therefore we create combinations of superpixels from the power set of $\mathds{P}_{i}$, denoted $2^{\mathds{P}_{i}}$.
@@
For an NN $\mathcal{N}$, an input $x$, and a partition $\mathds{P}_{i}$, we can generalize Def. 4.2 to the set of superpixels defined by $\mathds{P}_{i}$. We denote by $r_{i}(P_{i,j},x,\mathcal{N}(x))$ the degree of responsibility of a superpixel $P_{i,j}$ for $\mathcal{N}$’s classification of $x$, given $\mathds{P}_{i}$.
@@
We compute $r_{i}(P_{i,j},x,\mathcal{N}(x))$, the responsibility of each superpixel $P_{i,j}$ in the classification of $x$, in Algorithm 1. For a superpixel $P_{i,j}$, we define the set
@@
[…]
@@
OUTPUT: a responsibility map $\mathds{P}_{i}\rightarrow\mathds{Q}$
@@
5.2 Iterative refinement of responsibility
Algorithm 1 calculates the responsibility of each superpixel, subject to a given partition. Then, it proceeds with only the high-responsibility superpixels. Note that in general, it is possible that all superpixels in a given partition have the same responsibility. Consider, for example, a situation where the explanation is right in the middle of the image, and our partition divides the image into four quadrants. Each quadrant would be equally important for the classification, hence we would not gain any insight into why the image was classified in that particular way. In this case, the algorithm chooses another partition. It can also be the case that there might be multiple disjoint combinations of high responsibility superpixels in $\mathds{P}_{i}$. Refining them all can be computationally expensive. rex allows the user to choose from various pruning strategies to reduce the amount of work performed.
Different combinations of superpixels may have equal explanatory power. In the event that one superpixel is all that is required, therefore having responsibility 1, the further iterative subdivision of that superpixel is easy. However, if responsibility is split over multiple superpixels, we need a more refined approach. Take, for example, a situation where the right hand side of the image contains the explanation, as can be seen in Figure 6(b). Here responsibility is split over two super pixels. Let us say that the passing combination of superpixels $P_{c}=\{0,2\}$, assuming that we have numbered $P_{i,j}$ consecutively, and that $s=4$. Each superpixel $P_{i,j}$ has responsibility 0.50.50.5. We cannot simply mask superpixel 2 while the refining superpixel 0 as both together are required for the classification. We handle this problem by holding one superpixel to its original value while refining the other and then flipping the procedure. The superpixel held to its original value, but not being refined, does not get additional responsibility. This procedure is relatively scalable to sets containing more than 2 superpixels. In our implementation, rex, $s$ is hard-coded to 444 because this result in powerset of superpixels of manageable size and the refinement process for sets of superpixels just discussed needs to handle at most three superpixels at a time.
[…]

### v3

@@ 4. Theoretical Foundations of Our Approach
Unfortunately, the precise computation of explanations is intractable—see Appendix A for the proof. It is of little surprise, as all the existing definitions of explanations are intractable, except in some very simple special cases. There is, therefore, a justification for an approximation algorithm for explanations. Section 5, described in the next section, computes approximate explanations. It uses an approximate ranking of pixels according to their importance for the classification of the input image as a heuristic to guide explanation discovery. We formalize the notion of importance as follows.
Definition 4.2 (Sufficient responsibility).
The degree of sufficient responsibility of an explanation $\vec{V}_{exp}$ for the classification of $x$ by $\mathcal{N}$ is defined as $1/|\vec{V}_{exp}|$. We also extend this value to all pixels in $\vec{V}_{exp}$. That is, $v_{i}$ (and its matching pixel $p_{i}$) have the degree of sufficient responsibility $1/|\vec{V}_{exp}|$ for the classification of $x$ by $\mathcal{N}$, where $\vec{V}_{exp}$ is the smallest explanation for the classification of $x$ that contains $v_{i}$. If there is no explanation that contains $v_{i}$, then its degree of sufficient responsibility is defined as 0.
Similarly to the definition of explanation, Definition 4.2 has a match in the existing causality landscape, as we show in Appendix A, where we also prove its intractability. The following observation describes a brute-force approach to computing sufficient responsibility, which is clearly exponential in the number of pixels of the image.
Observation 4.3.
Given an image $x$ and its classification $o$, we can calculate the degree of sufficient responsibility of each pixel $p_{i}$ of $x$ by directly applying Definition 4.1, that is, by checking the conditions EXIM1 and EXIM2 for all subsets of pixels of $x$.
In the next section, we describe an efficient approximation algorithm for computing explanations that is based on an efficiently computable approximate degree of responsibility.
5. The rex Algorithm
@@
Figure 3. High level overview of Algorithm 3. The causal ranking algorithm produces an approximate responsibility map (①). The pixels in the image are then ordered by their approximate responsibility, and the explanation extraction algorithm uses this ranking to produce an approximately minimal sufficient explanation (②), which captures the information required for the DNN to give the classification ('starfish' in this example).
In this section we introduce our algorithm for computing approximate explanations. The high-level view of the algorithm is presented in Figure 3. The algorithm consists of two independent parts. The first part ranks pixels according to their approximate degree of sufficient responsibility (Definition 4.2); the second part is a greedy algorithm that extracts (approximate) explanations from this ranking.
The algorithm for computing an approximate degree of responsibility is based on the notion of a superpixel $P_{i}$. We slightly abuse the definition of a superpixel as used in lime, where it refers to pixels which are clustered via some segmentation algorithm. In our context, however, it is simply some subset of pixels of a given image which have not been masked. rex does not rely on any externally provided initial segmentation. The algorithm instead partitions an input image into a small number of rectilinear superpixels, computes their degree of sufficient responsibility for the output, and then iteratively refines the superpixels with responsibility exceeding some predefined threshold.
@@
The computation of the degree of sufficient responsibility of each superpixel for the output is precise and is described in Section 5.1. The approximation comes from the iterative refinement step in the algorithm, described in Section 5.2, and is due to the fact that the final ranking of pixels depends on the selected partition. To ameliorate the effect of a particular partition, the algorithm is repeated a number of times with partitions selected independently at random, and the results are averaged across all the partitions. Section 5.3 describes the averaging process and the greedy construction of an approximate sufficient explanation from the ranked list of pixels. Section 5.4 gives a step-by-step example to illustrate the working of the algorithm.
The scalability of the approach relies on the following observation.
Observation 5.1.
The pixels with the highest responsibility for the model’s decision are located in superpixels with the highest responsibility.
Intuitively, the observation holds when pixels with high responsibility do not appear in the superpixels surrounded by other pixels with very low responsibility for the input image classification outcome. Although this can happen in principle, especially in the case of adversarial images, we do not encounter this in practice. While the model has access to data only, this data has strong correlations due to the underlying data production mechanism.
5.1. Computing the degree of responsibility of a superpixel
@@
Given a set of pixels $\mathcal{P}$, we use $\mathds{P}_{i}$ to denote a partition of $\mathcal{P}$, that is, a set $\{P_{i,j}:\bigcup P_{i,j}=\mathcal{P}$ and $\forall j\not=k,P_{i,j}\cap P_{i,k}=\emptyset\}$. The number of elements in $\mathds{P}_{i}$ is a parameter, denoted by $s$; in our implementation, rex, we set $s=4$ as this provides convenient partitioning. We refer to $P_{i,j}$ as superpixels. It is insufficient to consider each superpixel in isolation: there is no reason why any one superpixel should be the cause of a classification. Therefore we create combinations of superpixels from the power set of $\mathds{P}_{i}$, denoted $2^{\mathds{P}_{i}}$.
@@
For an NN $\mathcal{N}$, an input $x$, and a partition $\mathds{P}_{i}$, we can generalize Definition 4.2 to the set of superpixels defined by $\mathds{P}_{i}$. We denote by $r_{i}(P_{i,j},x,\mathcal{N}(x))$ the sufficient responsibility of a superpixel $P_{i,j}$ for $\mathcal{N}$’s classification of $x$, given $\mathds{P}_{i}$.
@@
We compute $r_{i}(P_{i,j},x,\mathcal{N}(x))$, which is an approximation of the degree of sufficient responsibility (Definition 4.2) of each superpixel $P_{i,j}$ for the classification of $x$, in Algorithm 1. For a superpixel $P_{i,j}$, we define the set
@@
For a mutant image $x_{m}$, we define $\mathit{diff_{i}}(x_{m},x)$ as the number of superpixels in the partition $\mathds{P}_{i}$ that are masked in $x_{m}$ (that is, the difference between $x$ and $x_{m}$ with respect to $\mathds{P}_{i}$). For an image $y$, we denote by $y(P_{i,j})$ an image that is obtained by masking the superpixel $P_{i,j}$ in $y$. The responsibility of a superpixel $P_{i,j}$ is calculated by Algorithm 1 as a minimum difference between a mutant image and the original image over all mutant images $x_{m}$ that do not mask $P_{i,j}$, are classified the same as the original image $x$, and masking $P_{i,j}$ in $x_{m}$ changes the classification.
Algorithm 1 $\mathit{responsibility}(x,\mathds{P}_{i})$
@@
OUTPUT: a responsibility map $R:\mathds{P}_{i}\rightarrow\mathds{Q}$
@@
5.2. Iterative refinement of responsibility
Algorithm 1 calculates the responsibility of each superpixel, subject to a given partition. It then proceeds with only the high-responsibility superpixels. It returns a responsibility map, $R$, which maps superpixels to the rational numbers $\mathds{Q}$, indicating their degree of responsibility.
Note that in general, it is possible that all superpixels in a given partition have the same responsibility. Consider, for example, a situation where the explanation is right in the middle of the image, and our partition divides the image into four quadrants. Each quadrant would be equally important for the classification, hence we would not gain any insight into why the image was classified in that particular way. In this case, the algorithm starts again from another partition. It can also be the case that there might be multiple disjoint combinations of high responsibility superpixels in $\mathds{P}_{i}$. Refining them all can be computationally expensive. rex allows the user to choose from various pruning strategies to reduce the amount of work performed.
Different combinations of superpixels may have equal explanatory power. In the event that one superpixel is all that is required, therefore having responsibility 1, the further iterative subdivision of that superpixel is easy. However, if responsibility is split over multiple superpixels, we need a more refined approach. Take, for example, a situation where the right hand side of the image contains the explanation, as can be seen in Figure 10(b). Here responsibility is split over two super pixels. Let us say that the passing combination of superpixels $P_{c}=\{0,2\}$, assuming that we have numbered $P_{i,j}$ consecutively, and that $s=4$. Each superpixel $P_{i,j}$ has responsibility $0.5$. We cannot simply mask superpixel $2$ while refining superpixel 0 as both together are required for the classification. We handle this problem by holding one superpixel to its original value while refining the other and then reversing the procedure. The superpixel held to its original value, but not being refined, does not get any additional responsibility.
Algorithm 2 $\mathit{iterative\_responsibility\_refinement}(x,\mathds{P}_{i})$